# Multi-Fidelity Bayesian Optimization for Nash Equilibria with Black-Box Utilities

### v1

@@ 14             Set episode time $\tau=\tau+1$
III-D Evaluation Phase
$\hskip 14.22636pt+\beta_{n,j}\sigma^{(M)}_{n,j,t_{j}}(\mathbf{x})],$ (31)
@@ IV Theoretical Analysis
In this section, we extend the proof of the sublinearity of the regret (15) in the general case with multiple fidelity levels and players.
@@ IV-A Technical Assumptions
We start by separating the contribution of different episodes $\{\mathcal{E}_{j}\}_{j=1}^{J}$ to the regret $R(\Lambda)$ in (15). This is done by defining the episode regret $R(\mathcal{E}_{j})$. As detailed next, this captures the gap between the accumulated loss (12) incurred within the episode $\mathcal{E}_{j}$ and the optimal dissatisfaction performance that can be achieved given the episode budget $\Lambda_{j}$, namely $(\Lambda_{j}/N\lambda^{(M)})\epsilon^{*}$.
@@
$\Lambda_{j}=\sum_{n\in\mathcal{N}}\sum_{\tau=1}^{t_{j}}\lambda^{(m_{n,j,\tau})}+N\cdot\lambda^{(M)},$ (38)
@@
$R(\mathcal{E}_{j})=(t_{j}\cdot C+\epsilon_{j})-\frac{\Lambda_{j}}{N\lambda^{(M)}}\cdot\epsilon^{*},$ (40)
@@
for some real-valued coefficients $\{a_{s}\}_{s=1}^{S_{n}}$ and inducing vectors $\{\mathbf{x}_{s}\}_{s=1}^{S_{n}}$, where the number of terms in the sum, $S_{n}$, may be unbounded.
We also make the standard assumption that the complexity of each utility function $u_{n}(\mathbf{x})$, as measured by the RKHS norm $||u_{n}||_{\mathcal{H}_{k}}$, is bounded [34, 35].
@@
The RKHS norm of the utility function $u_{n}(\mathbf{x})$ satisfies the inequality $||u_{n}||_{\mathcal{H}_{k}}\leq B$, for all players $n\in\mathcal{N}$ for some known upper bound $B>0$.
Figure 4: Visualization of MF-UCB-PNE optimization trajectories against a heat map of each player’s utility function. The total query budget is set to $\Lambda=32$, and the threshold in (29) is set to $\eta=0.5$. The indexes $(t,m_{n,t})$ associated to each point in the heat map describe the time step $t$ increasing the counter across episodes and the fidelity decision $m_{n,t}$ assigned to player $n$ at time step $t$. The $\epsilon^{*}$-PNE is marked as a red star. The initial action at time $t=1$ and the actions made in evaluation phases are shown as red squares, while decisions made in the exploration phases are marked as black circles. Dashed black arrows follow the updates across time steps, while solid red arrows connect the actions obtained at the evaluation phases at the end of each episode.
@@ IV-B Regret Analysis
Lemma 1 (Uniform Error Bound in RKHS [34, Theorem 2]).
@@
Lemma 2 (Dissatisfaction Coverage [28, Lemma 1]).
@@
are simultaneously valid for all inputs $\mathbf{x}\in\mathcal{X}$ and episodes $j\in\mathcal{J}$.
@@
$R(\Lambda)\leq\bigg{(}\frac{C}{\lambda^{(1)}}-\frac{\epsilon^{*}}{\lambda^{(M)}}\bigg{)}\sqrt{\Lambda}\cdot\gamma_{J}+4\beta_{J}\sqrt{4(J+2)\gamma_{J}},$ (48)
@@
Leveraging the per-episode decomposition (40), the regret upper bound (48) is the sum of a regret term incurred while exploring lower fidelity utilities and of a regret term that accounts for the dissatisfaction levels accrued during the evaluation phases. Specifically, the first term in (48) is derived by considering the largest regret obtained during the exploration phases by selecting the minimum fidelity levels $m_{n,j,\tau}=1$ for all players $n$ and episodes $j$. In contrast, the second term in (48) is akin to the regret bound for the single-fidelity multi-player case studied in [28].
@@ V-B Synthetic Random Utility Functions
For the first experiment, we consider a game with $N=2$ players and $M=2$ fidelity levels. The $2\times 1$ multi-fidelity function vector $\mathbf{u}_{n}(\mathbf{x})$ for each player $n$ is sampled from the MOGP prior with lengthscale parameters in the RBF kernels (20) and (21) given by $h=0.89$ and $\zeta^{(1)}=\zeta^{(2)}=0.78$, respectively. The correlation parameter in the auto-regressive model (19) is set as $\rho^{(1)}=0.768$, and the observations noise variance is $\sigma^{2}=0.1$. The action space for each player is restricted to the interval $\mathcal{X}_{n}=[-1,1]$, and the cost levels are $\lambda^{(1)}=1$ and $\lambda^{(2)}=8$. For all schemes, we consider choices both a well-specified MOGP, which adopts the same parameters used in the random functions generation process, and a misspecified MOGP, which sets different parameters $h=0.62,\zeta^{(1)}=\zeta^{(2)}=0.41$ and $\rho^{(1)}=0.625$.
@@ V-C Optimizing Power Allocation in Wireless Systems
The centralized optimizer can estimate the utilities (50) by observing $\lambda$ realizations of the random variable $\log(1+\text{SINR}_{n})$. The cost is given by the number of realizations, and we allow for the options $\lambda^{(1)}=1,\lambda^{(2)}=10,\lambda^{(3)}=20,\lambda^{(4)}=50$ and $\lambda^{(5)}=100$. We assume $N=20$ links, each with normalized transmission power $-13$ dB $\leq x_{n}\leq 23$ dB, and noise power $\sigma_{z}^{2}=-20$ dB. We simulate Rayleigh fading direct channels $h_{n,n}\sim\mathcal{CN}(0,1)$ and interference channels $h_{n,m}\sim\mathcal{CN}(0,\psi)$ with $\psi=-20$ dB. The MOGP is initialized with parameters $h=0.89,\zeta^{(m)}=0.78$, and $\rho=0.768$. Note that the MOGP is not well-specified, since data generation does not follow an MOGP. For MF-UCB-PNE, we consider a fixed threshold $\eta=0.4$, as well as the optimal $\eta$ obtained in a similar manner to Fig. 6.
@@ V-D Optimizing a Random Access Protocol
We consider a scenario of $N=5$ mobile terminals with energy consumptions $c_{1}=50$ units for staying active and $c_{2}=70$ units for transmission, while the maximum energy consumption constraints for the mobile terminals are set to $\{60,55,50,45,40\}$. The MOGP is initialized with parameters $h=1.08,\zeta^{(m)}=0.41,\rho=0.797$ for all players. And we set a threshold $\eta=0.2$ as well as the optimal $\eta$ for MF-UCB-PNE. We set $\xi_{n}=6.5\times 10^{-4}$ for all players, and low-fidelity evaluations of the utility function use the factors $\omega^{(1)}=4.9\times 10^{-4},\omega^{(2)}=5.5\times 10^{-4}$, and $\omega^{(3)}=6.1\times 10^{-4}$ instead, with the corresponding query costs $\lambda^{(1)}=1,\lambda^{(2)}=5,\lambda^{(3)}=10$ and $\lambda^{(4)}=20$.
@@ VI Conclusion
The paper has introduced a centralized optimization framework for identifying approximate pure Nash equilibria (PNE) in complex multi-agent systems with black-box utility functions. The framework was devised within a Stackelberg game formulation in which a central controller coordinates multiple strategic players. To tackle the challenge of learning unknown utilities through sequential queries, the paper presented MF-UCB-PNE, a novel multi-fidelity Bayesian optimization algorithm. MF-UCB-PNE explicitly balances exploration and exploitation across fidelity levels under stringent query cost constraints. By leveraging inexpensive low-fidelity evaluations for broad exploration and transitioning to high-fidelity queries for refined exploitation, the algorithm efficiently converges toward configurations that approximate PNE solutions while respecting budget limits. The paper provides a rigorous theoretical regret analysis, proving that MF-UCB-PNE achieves asymptotic no-regret performance under suitable conditions. Empirically, results show that the proposed approach consistently identifies high-quality equilibrium configurations within limited query budgets.
@@
$R(\mathcal{E}_{j})$ $=(t_{j}\cdot C+\epsilon_{j})-\frac{\Lambda_{j}}{N\lambda^{(M)}}\epsilon^{*}$
$\leq\frac{\Lambda_{j}-N\lambda^{(M)}}{N\lambda^{(1)}}C+\epsilon_{j}-\frac{\Lambda_{j}}{N\lambda^{(M)}}\epsilon^{*},$ (53)
where we set $t_{j}=\frac{\Lambda_{j}-N\lambda^{(M)}}{N\lambda^{(1)}}$ as the maximum number of steps during the exploration phase in episode $\mathcal{E}_{j}$. Assuming that MF-UCB-PNE terminates within $J$ episodes, by Definition 2, the regret upper bound can be then obtained as
$R(\Lambda)$ $\leq\frac{\Lambda-JN\lambda^{(M)}}{N\lambda^{(1)}}C+\sum_{j=1}^{J}\epsilon_{j}-\frac{\Lambda}{N\lambda^{(M)}}\epsilon^{*}$ (54a)
$=\frac{\Lambda-JN\lambda^{(M)}}{N\lambda^{(1)}}C-\frac{\Lambda-JN\lambda^{(M)}}{N\lambda^{(M)}}\epsilon^{*}+\sum_{j=1}^{J}(\epsilon_{j}-\epsilon^{*})$ (54b)
$=\bigg{(}\frac{C}{N\lambda^{(1)}}-\frac{\epsilon^{*}}{N\lambda^{(M)}}\bigg{)}\sum_{j=1}^{J}\Lambda_{\mathcal{L}_{j}}+\sum_{j=1}^{J}(\epsilon_{j}-\epsilon^{*}).$ (54c)
The inequality (54a) is obtained by summing the episode regret upper bound in (53); equation (54b) subtracts the best achievable dissatisfaction $J\epsilon^{*}$ from (54a) across $J$ episodes, and applies the associative law. The first term in (54c) represents the regret incurred from decision sequences in exploration phases, while the second term in (54c) describes the regret obtained by optimizing the utility functions at the maximum fidelity level.
@@
where the inequality (57b) is obtained by introducing the upper and lower confidence intervals obtained on the selected action profile $\mathbf{x}_{j,t_{j}+1}$ instead of the $\epsilon^{*}$-PNE solution $\mathbf{x}^{*}$; (57c) is obtained by aligning the player index which has the maximum dissatisfaction upper confidence bound as in (34); (57d) is obtained by applying the union bound to Lemma 2; and the inequality (57f) is obtained from [34, Lemma 4].
@@
When the fidelity decision at each time step is selected as $M$, we only perform the evaluation phase. For brevity of notation, we omit the episode index and let $\mathbf{M}_{n,t-1}=M\cdot\mathbf{1}$ denote the fidelity decisions for all the previous $t-1$ time steps in the optimization process. The maximal mutual information in (58) can be then inductively expressed as
[…]
@@
which indicates that the maximal mutual information (59) is obtained by always querying the action profiles at the maximum-fidelity level $M$. Since the maximal information gain $\gamma_{n,t}$ satisfies submodularity, a diminishing returns property [42], it can be upper bounded via a greedy procedure at each time [16, Section V.A], i.e.,
@@
Given the selected RBF kernel in MOGP, the maximal information gain is upper bounded by $\gamma_{n,t}=\mathcal{O}(\log(t)^{d+1})$ [16, Theorem 5]
@@
$R(\Lambda)$
$\leq\bigg{(}\frac{C}{\lambda^{(1)}}-\frac{\epsilon^{*}}{\lambda^{(M)}}\bigg{)}\sqrt{\Lambda}\cdot\gamma_{J}+4\beta_{J}\sqrt{4(J+2)\gamma_{J}}$ (61a)
$\leq\mathcal{O}(\sqrt{\Lambda}\log(T_{J})^{d+1})+\mathcal{O}(\sqrt{J}\log(T_{J})^{d+1})$ (61c)
$\leq\mathcal{O}\bigg{(}(\sqrt{J}+\sqrt{\Lambda})\log\Big{(}\frac{\Lambda-J\lambda^{(M)}}{\lambda^{(1)}}+J\Big{)}^{d+1}\bigg{)},$ (61e)
where inequality (61c) is obtained by using $\beta_{J}=\mathcal{O}(\sqrt{\gamma_{J}})$, and applying the maximal information gain upper bound; and inequality (61e) is obtained by considering the maximum number of steps one can query given budget $\Lambda$. The upper bound (61e) satisfies the limit (37), hence, MF-UCB-PNE incurs asymptotic no-regret performance.

### v2

@@ 14      Set episode time $\tau=\tau+1$
Like BO, MF-UCB-PNE is a meta-algorithm that can be instantiated using different implementations for the update of the GP surrogate and for the acquisition functions (29) and (34). For context, in Bayesian optimization, surrogate model complexity ranges from $\mathcal{O}(t^{3})$ for exact GP inference to linear complexity when using sparse GPs with inducing points. Similarly, acquisition function optimization complexity varies with the chosen approach [44]. Our framework inherits this flexibility and can be implemented with the same range of computational strategies.
The key distinction from standard Bayesian optimization is that our exploitation phase requires solving for a Nash equilibrium. While this is computationally challenging (as is the optimization problem in standard BO), numerous efficient heuristic solutions exist and can be readily employed. This was comprehensively explained in our response. Finding a Nash equilibrium is, in general, computationally intractable in worst-case scenarios, even for simple two-player games. This inherent hardness means that our algorithm, like [40], can be implemented exactly in settings in which the bilevel optimization problem (34) can be solved via convex optimization or exhaustive search [36, 16, 2], while otherwise requiring the use of approximation methods [37].
@@ IV Theoretical Analysis
In this section, we extend the proof of the sublinearity of the regret (16) in general case with multiple fidelities and players.
@@ IV-A Technical Assumptions
We start by separating the contribution of different episodes $\{\mathcal{E}_{j}\}_{j=1}^{J}$ to the regret $R(\Lambda)$ in (16). This is done by defining the episode regret $R(\mathcal{E}_{j})$. As detailed next, this captures the gap between the accumulated reward (13) incurred within the episode $\mathcal{E}_{j}$ and the optimal dissatisfaction performance that can be achieved given the episode budget $\Lambda_{j}$, namely $(1-\epsilon^{*}/C)(\Lambda_{j}/N)$.
@@
$\Lambda_{j}=\sum_{n\in\mathcal{N}}\sum_{\tau=1}^{t_{j}}\lambda^{(m_{n,j,\tau})}+N,$ (39)
@@
$R(\mathcal{E}_{j})=\frac{\Lambda_{j}}{N}\Big(1-\frac{\epsilon^{*}}{C}\Big)-\Big(1-\frac{\epsilon_{j}}{C}\Big),$ (41)
@@
for some real-valued coefficients $\{a_{s}\}_{s=1}^{S_{n}}$ and inducing vectors $\{\mathbf{x}_{s}\}_{s=1}^{S_{n}}$, where $S_{n}$ may be unbounded.
We also make the standard assumption that the complexity of each utility function $u_{n}(\mathbf{x})$, as measured by the RKHS norm $u_{n}||_{\mathcal{H}_{k}}$, is bounded [6, 49].
@@
The RKHS norm of the utility function $u_{n}(\mathbf{x})$ satisfies the inequality $u_{n}||_{\mathcal{H}_{k}}\leq B$, for all players $n\in\mathcal{N}$ for some known upper bound $B>0$.
Figure 4: Visualization of MF-UCB-PNE optimization trajectories against a heatmap of each player’s utility function. The query budget is set to $\Lambda=32$, and the threshold in (30) is set to $\eta=0.5$. The indexes $(t,m_{n,t})$ associated to each point in the heat map describe the time step $t$ increasing the counter across episodes and the fidelity decision $m_{n,t}$ assigned to player $n$ at time step $t$. The $\epsilon^{*}$-PNE is marked as a red star. The initial action at time $t=1$ and the actions made in evaluation phases are shown as red squares, while decisions made in the exploration phases are marked as black circles. Dashed black arrows follow the updates across time steps, while solid red arrows connect the actions obtained at the evaluation phases at the end of each episode.
@@ IV-B Regret Analysis
Lemma 1 (Uniform Error Bound in RKHS [6, Theorem 2]).
@@
Lemma 2 (Dissatisfaction Coverage [40, Lemma 1]).
@@
are simultaneously valid for all $\mathbf{x}\in\mathcal{X}$ and episodes $j\in\mathcal{J}$.
@@
$R(\Lambda)\leq\bigg(1-\frac{\epsilon^{*}}{C}\bigg)\sqrt{\Lambda}\cdot\gamma_{J}+\frac{4\beta_{J}}{C}\sqrt{4(J+2)\gamma_{J}},$ (49)
@@
Leveraging the per-episode decomposition (41), the regret upper bound (49) is the sum of a regret term incurred while exploring lower fidelity utilities and of a regret term that accounts for the dissatisfaction levels accrued during the evaluation phases. Specifically, the first term in (49) is derived by considering the largest regret obtained during the exploration phases. In contrast, the second term in (49) is akin to the regret bound for the single-fidelity multi-player case studied in [40].
@@ V-B Synthetic Random Utility Functions
For the first experiment, we consider a game with $N=2$ players and $M=2$ fidelity levels. The $2\times 1$ multi-fidelity function vector $\mathbf{u}_{n}(\mathbf{x})$ for each player $n$ is sampled from the MOGP prior with lengthscale parameters in the RBF kernels (21) and (22) given by $h=0.89$ and $\zeta^{(1)}=\zeta^{(2)}=0.78$, respectively. The correlation parameter in the auto-regressive model (20) is set as $\rho^{(1)}=0.768$, and the observations noise variance is $\sigma^{2}=0.1$. The action space for each player is restricted to the interval $\mathcal{X}_{n}=[-1,1]$, and the cost levels are $\lambda^{(1)}=1$ and $\lambda^{(2)}=8$. For all schemes, we consider choices both a well-specified MOGP, which adopts the same parameters used in the random functions generation process, and a misspecified MOGP, which sets different parameters $h=0.62,\zeta^{(1)}=\zeta^{(2)}=0.41$ and $\rho^{(1)}=0.625$. To evaluate the $\epsilon$-PNE of the synthetic game, we uniformly discretize the continuous action space $\mathcal{X}_{n}$ into $128$ strategies for every player, and we find a PNE by a brute-force grid search.
@@ V-C Optimizing Power Allocation in Wireless Systems
The centralized optimizer can estimate the utilities (51) by observing $\lambda$ realizations of the random variable $\log(1+\text{SINR}_{n})$. The cost is given by the number of realizations, and we allow for the options $\lambda^{(1)}=1,\lambda^{(2)}=10,\lambda^{(3)}=20,\lambda^{(4)}=50$ and $\lambda^{(5)}=100$. We assume $N=20$ links, each with normalized transmission power 13-13 dB $\leq x_{n}\leq 23$ dB, and noise power $\sigma_{z}^{2}=-20$ dB. We simulate Rayleigh fading direct channels $h_{n,n}\sim\mathcal{CN}(0,1)$ and interference channels $h_{n,m}\sim\mathcal{CN}(0,\psi)$ with $\psi=-20$ dB. The MOGP is initialized with parameters $h=0.89,\zeta^{(m)}=0.78$, and $\rho=0.768$. Note that the MOGP is not well-specified, since data generation does not follow an MOGP. For MF-UCB-PNE, we consider a fixed threshold $\eta=0.4$, as well as the optimal $\eta$ obtained in a similar manner to Fig. 6.
@@ V-D Optimizing a Random Access Protocol
We consider a scenario of $N=5$ mobile terminals with energy consumptions $c_{1}=50$ units for staying active and $c_{2}=70$ units for transmission, while the maximum energy consumption constraints for the mobile terminals are set to $\{60,55,50,45,40\}$. The MOGP is initialized with parameters $h=1.08,\zeta^{(m)}=0.41,\rho=0.797$ for all players. And we set a threshold $\eta=0.2$ as well as the optimal $\eta$ for MF-UCB-PNE. We set $\xi_{n}=6.5\times 10^{-4}$ for all players, and low-fidelity evaluations of the utility function use the factors $\omega^{(1)}=4.9\times 10^{-4},\omega^{(2)}=5.5\times 10^{-4}$, and $\omega^{(3)}=6.1\times 10^{-4}$ instead, with the corresponding query costs $\lambda^{(1)}=1,\lambda^{(2)}=5,\lambda^{(3)}=10$ and $\lambda^{(4)}=20$. The equilibrium solution is computed by fixed-point iterations [42].
@@
Moreover, MF-UCB-PNE exhibits an earlier and steeper decline in regret over low to moderate budgets, indicating that lower fidelity exploration has already localized promising regions before substantial high-fidelity queries with high costs. Near the terminal budget, MF-UCB-PNE shows a final sharp contraction as the remaining budget is concentrated on high-fidelity checks around the best solutions obtained so far, whereas UCB-PNE and PE tend to flatten after uniform high-fidelity sampling. This early localization trajectory followed by targeted certification aligns with the two-phase design and explains the lower regret achieved under the same total budget.
@@ VI Conclusion
The paper has introduced a centralized optimization framework for identifying approximate pure Nash equilibria (PNE) in complex multi-agent systems with black-box utility functions. The framework was devised within a learned optimization problem with an equilibrium constraint in which a central controller coordinates multiple strategic players. To tackle the challenge of learning unknown utilities through sequential queries, the paper presented MF-UCB-PNE, a novel multi-fidelity Bayesian optimization algorithm. MF-UCB-PNE explicitly balances exploration and exploitation across fidelity levels under stringent query cost constraints. By leveraging inexpensive low-fidelity evaluations for broad exploration and transitioning to high-fidelity queries for refined exploitation, the algorithm efficiently converges toward configurations that approximate PNE solutions while respecting budget limits. The paper provides a rigorous theoretical regret analysis, proving that MF-UCB-PNE achieves asymptotic no-regret performance under suitable conditions. Empirically, results show that the proposed approach consistently identifies high-quality equilibrium configurations within limited query budgets.
@@
$R(\mathcal{E}_{j})=\frac{\Lambda_{j}}{N}\Big(1-\frac{\epsilon^{*}}{C}\Big)-\Big(1-\frac{\epsilon_{j}}{C}\Big).$ (54)
Assuming that MF-UCB-PNE terminates within $J$ episodes, by Definition 2, the regret upper bound can be then obtained as
$R(\Lambda)$ $=\frac{\Lambda}{N}\Big(1-\frac{\epsilon^{*}}{C}\Big)-\sum_{j=1}^{J}\Big(1-\frac{\epsilon_{j}}{C}\Big)$ (55a)
$=\Big(\frac{\Lambda}{N}-J\Big)\Big(1-\frac{\epsilon^{*}}{C}\Big)+\sum_{j=1}^{J}\frac{\epsilon_{j}-\epsilon^{*}}{C}$ (55b)
$=\Big(\frac{1-\epsilon^{*}/C}{N}\Big)\sum_{j=1}^{J}\Lambda_{\mathcal{L}_{j}}+\sum_{j=1}^{J}\frac{\epsilon_{j}-\epsilon^{*}}{C}.$ (55c)
The inequality (55a) is obtained by summing the episode regret in (54); equation (55b) subtracts the best achievable dissatisfaction $J\epsilon^{*}$ from (55a) across $J$ episodes, and applies the associative law. The first term in (55c) represents the regret incurred from decision sequences in exploration phases, while the second term in (55c) describes the regret obtained by optimizing the utility functions at the maximum fidelity level.
@@
where the inequality (58b) is obtained by introducing the upper and lower confidence intervals obtained on the selected action profile $\mathbf{x}_{j,t_{j}+1}$ instead of the $\epsilon^{*}$-PNE solution $\mathbf{x}^{*}$; (58c) is obtained by aligning the player index which has the maximum dissatisfaction upper confidence bound as in (35); (58d) is obtained by applying the union bound to Lemma 2; and the inequality (58f) is obtained from [6, Lemma 4].
@@
When the fidelity decision at each time step is selected as $M$, we only perform the evaluation phase. For brevity of notation, we omit the episode index and let $\mathbf{M}_{n,t-1}=M\cdot\mathbf{1}$ denote the fidelity decisions for all the previous t1t-1 time steps in the optimization process. The maximal mutual information in (59) can be then inductively expressed as
$\max\limits_{\begin{subarray}{c}\mathbf{x}\in\mathcal{X}\\ m_{n}\in\mathcal{M}\end{subarray}}\Bigg\{$ $\mathrm{I}(y_{n}^{(m_{n})};u_{n}(\mathbf{x})|\mathbf{x},m_{n},\mathbf{X}_{t-1},\mathbf{M}_{n,t-1})$
@@
which indicates that the maximal mutual information (60) is obtained by always querying the action profiles at the maximum-fidelity level $M$. Since the maximal information gain $\gamma_{n,t}$ satisfies submodularity, a diminishing returns property [22], it can be upper bounded via a greedy procedure at each time [39, Section V.A], i.e.,
@@
Given the selected RBF kernel in MOGP, the maximal information gain is upper bounded by $\gamma_{n,t}=\mathcal{O}(\log(t)^{d+1})$ [39, Theorem 5]
@@
$R(\Lambda)$ $\leq\bigg(1-\frac{\epsilon^{*}}{C}\bigg)\sqrt{\Lambda}\cdot\gamma_{J}+\frac{4\beta_{J}}{C}\sqrt{4(J+2)\gamma_{J}}$ (62a)
$\leq\mathcal{O}(\sqrt{\Lambda}\log(T_{J})^{d+1})+\mathcal{O}(\sqrt{J}\log(T_{J})^{d+1})$ (62b)
$\leq\mathcal{O}\bigg((\sqrt{J}+\sqrt{\Lambda})\log\Big(\frac{\Lambda-J\lambda^{(M)}}{\lambda^{(1)}}+J\Big)^{d+1}\bigg),$ (62c)
where inequality (62b) is obtained by using $\beta_{J}=\mathcal{O}(\sqrt{\gamma_{J}})$, and applying the maximal information gain upper bound; and inequality (62c) is obtained by considering the maximum number of steps one can query given budget $\Lambda$. The upper bound (62c) satisfies the limit (38), hence, MF-UCB-PNE incurs asymptotic no-regret performance.